OpenStudy (cchavez):

Please help... Which equation is the equation of a line that passes through (–10, 3) and is perpendicular to y = 5x – 7?

OpenStudy (will.h):

perpendicular to y = 5x – 7 means that the other slope will be the reciprocal Thus the slope is -1/5 passes through (-10,3) use the formula y - y1 = m(x - x1) where m is the slope y - 3 = -1/5(x - (-10) y - 3 = -1/5 (x + 10) y - 3 = (-1/5)x -2 y = (-1/5)x + 1

OpenStudy (cchavez):

a. y = 5x + 53 b. y = –The term shows 1 over 5.x – 7 c. y = –The term shows 1 over 5.x + 1 d. y = The term shows 1 over 5.x + 5

OpenStudy (cchavez):

c then. Thank you so much for showing how you got the answer.

OpenStudy (cchavez):

I am not good in this stuff at all

OpenStudy (will.h):

i answered it even before u type the answers wow lol

OpenStudy (will.h):

and you welcome let me know if you have other questions

OpenStudy (cchavez):

ok thanks again